Consider a market defined as follows: Demand: Q = 400 – P Supply: Q = 2P – 200 Additionally, a negative externality of $75 per unit is associated with the traded good.

How many units should be traded in this market to maximize benefit to society?

If the government does not implement the policy you proposed in part (b), what is the size of the deadweight loss?
Answer #1

Answer

The negative externality is internalized to produce at the socially optimum level by taxing the market.

Adding the tax to the supply curve by converting the supply curve to inverse supply

Q=2P-200

2P=Q+200

P=0.5Q+100+75=0.5Q+175

converting back

0.5Q=P-175

Q=2P-350

equating to demand curve to find the equilibrium

400-P=2P-350

3P=750

P=250

Q=400-250=150 units

150 units should be traded in this market to maximize the benefit to society

===

Without tax, the market is in market equilibrium at Qd=Qs

400-P=2P-200

3P=600

P=200

Q=400-200=200

===

DWL=0.5*externality per unit * change in output

=0.5*75*(200-150)

=1875

The DWL is $1875